# SADMoE: Exploiting Activation Sparsity with Dynamic-$k$ Gating

### v2

@@ Appendix A Difference between router training in SADMoE and MoEfication
In comparison to their approach, the router training procedure in SADMoE differs in multiple aspects:
@@
We find that the above differences are responsible for the improved performance of our router (see Figure 8).
Appendix B Comparision of FLOPs between standard FFN layer and dynamic-$k$ MoE
@@
which shows that dynamic-$k$ expert selection only exceeds the FLOPs cost of the standard network when the dynamic-$k$ rule selects almost all experts or the number of experts becomes very high. For an even more detailed analysis, we refer to Figure 14 where the the cost ratio between our method and standard MLP is shown, assuming different router sizes and $e=4$ as standard for most Transformer models. In practice, we use $d_{h}=128$, so in all our experiments $d_{m}=6d_{h}$.
Appendix C Base model activation sparsity
In this section, we investigate the impact of activation sparsity levels on the performance of our method. We train multiple models with different activation sparsity enforcement loss weights $\alpha$ and convert them to Mixture-of-Experts with our method and MoEfication. Note that, in contrast to other sections in this work, we apply MoEfication to the model fine-tuned with sparsity enforcement.
The results highlight the positive correlation between the activation sparsity and the performance of the converted MoE, as higher sparsity in the base model always translates to better performance for both SADMoE and MoEfication. Our method achieves better results for every base model in all cases, which proves that dynamic-$k$ selection makes better use of the induced sparsity.
Figure 12: FLOPs vs loss for GPT-2 on OpenWebText for SADMoE (left) and MoEfication (right) with different sparsity levels in the base model. Sparsity enforcement is beneficial for both methods, but our one outperforms MoEfication baseline as dynamic-$k$ can better utilize the sparsity.
Appendix D Additional results with expert size and GELU
Figure 13: Impact of expert division granularity on our method with the GELU activation function.
In addition to experiments in Section 5.3, we present the results of similar ablation carried on the sparsified GPT-2 model with GELU activation. The results, presented in Figure 13, follow the same pattern as before, which supports our claim that the sparsification enables the GELU-based models to function similarly to ReLU-based ones.
Appendix E Expert activation patterns for attention projection layers
Following the analysis for MoE-converted FFN layers in Section 5.4, we investigate the activation patterns in MHA projections modified with our method in Figures 15, 16, 17 and 18. The projection modules display lower levels of sparsity than FFNs, which is to be expected as our projection layers have lower intermediate dimensionality. Expert selection distribution patterns in $Q$ and $K$ show significant similarity, and the patterns in $V$ and output projections are also similar to a lesser degree. The variance in the number of selected experts in MHA projections is higher than in FFN layers, but it still exists and the distribution in some of the layers seems to be bimodal, which still justifies dynamic-$k$ selection rule.
Appendix F Training details
In this Section, we describe the technical details used in the SADMoE conversion procedure. For full reproducibility, we also share the source code that we used for conducting the experiments. All experiments were performed using the PyTorch library (Paszke et al., 2019) on the NVIDIA A100 and V100 GPUs.
F.1 Language modeling
We base our code and hyperparameters on the nanoGPT repository provided at https://github.com/karpathy/nanoGPT. In all pretraining experiments, we keep the fixed number of steps at 1000 and keep the effective batch size equal to the value in the repository through gradient accumulation. The alpha values for sparsity enforcement can be found at Figure 12. We train the routers for SADMoE and MoEfication for 2000 steps using one GPU and tuning the learning rates for a given expert size from the range between $0.002-0.005$. For router training, we use Adam optimizer and cosine warmup scheduler.
F.2 Text classification
All experiments start from the same pre-trained BERT-base checkpoint. For methods requiring ReLU activation function, we replaced GELU with ReLU and continue model pretraining on concatenated wikipedia (Foundation, ) and books (Zhu et al., 2015) corpora for $5000$ steps on $8$ GPUs using main setup from https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_mlm.py, per device batch size $96$ and learning rate $5\cdot 10^{-4}$. For MHA projections replacement we use the same corpus and train replaced MLP modules on a single GPU with batch size $128$ and learning rate $0.001$ for $3000$ steps. We finetuned base dense models on CARER datasetw for $5$ epochs with $2\cdot 10^{-5}$ learning rate. For sparsity enforcement in SADMoE we use $\alpha$ linearly increasing from zero to $0.001$ over training. For both MoEfication and SADMoE we train routers with batch size 64 and initial learning rate $0.001$ for 5 epochs. In all experiments, we use Adam optimizer with linear learning rate decay. For MoEfication we use same expert size $32$, for SADMoE we use more granular expert size equals $6$. For ZTW we trained ICs for 5 epochs with batch size $32$ and learning rate $0.01$.
F.3 Image classification
All methods start with the same pre-trained ViT-B from the torchvision library and are trained on ImageNet-1k using the augmentation proposed by Touvron et al. (2022). We use mixup ($0.8$), cutmix, label smoothing ($0.1$), gradient clipping ($1.0$) and the Adam optimizer with cosine learning rate schedule without warm-up. For SADMoE, we replace the MHA projections and train the replacements for $3$ epochs with the initial learning rate $0.001$ and batch size $128$, and then finetune the model for $90$ epochs with sparsity enforcement weight $\alpha=0.2$, initial learning rate $2\cdot 10^{-5}$ and batch size $512$. We then convert the modules into MoE layers, and train the gating networks for $7$ epochs with the initial learning rate set to $0.001$ and batch size $128$. We train ZTW for $100$ epochs in total, allocating $5$ epochs for ensemble training, while keeping the rest of the original hyperparameters unchanged. For MoEfication, we first convert the pre-trained model to ReLU-based one and finetune for $90$ epochs with an initial learning rate of $0.0001$ and batch size $256$. We then split the weights and train the routers for $10$ epochs with the initial learning rate $0.001$ and batch size $256$.
Appendix G Contributions
Filip integrated GPT-2 into the codebase and ran GPT-2 experiments, performed the distribution of non-zero activations analysis, the expert granularity analysis, and the activation function analysis. He contributed to the design of dynamic-k gating, and played a primary role in designing the experiments, writing of the article and producing its visualizations.
Bartosz set the research direction of the project and proposed the alternative routing scheme, dynamic-k expert selection, the additional activation sparsity enforcement phase for ReLU and GELU, and the replacement of MHA projection layers. He also wrote the shared codebase for the experiments, carried out the ViT-B experiments, and also played a primary role in the writing and editing of the article.
Mikołaj made this paper possible by performing all of the experiments at the initial stages of the project and implementing MoEfication and numerous variants of our method. He carried out the BERT experiments, performed the expert selection pattern analysis and the ablation study, and contributed to the crafting of the paper.
Kamil helped with the writing and editing of the paper.

### v4

@@ Appendix A Difference between router training in D2DMoE and MoEfication
In comparison to their approach, the router training procedure in D2DMoE differs in multiple aspects:
@@
We find that the above differences are responsible for the improved performance of our router (see Figure 9(a)).
Appendix B Comparison of FLOPs between standard FFN layer and dynamic-$k$ MoE
@@
which shows that dynamic-$k$ expert selection only exceeds the FLOPs cost of the standard network when the dynamic-$k$ rule selects almost all experts or the number of experts becomes very high. For an even more detailed analysis, we refer to Figure 10 where the cost ratio between our method and standard MLP is shown, assuming different router sizes and $e=4$ as standard for most Transformer models. In practice, we use $d_{h}=128$, so in all our experiments $d_{m}=6d_{h}$.
@@
Appendix C Efficient implementation of D2DMoE
Figure 11: Wall-clock time measurements of the ViT-B model and its corresponding D2DMoE model.
In Listing 1 we present the pseudocode for our efficient implementation of the forward pass of the D2DMoE module. We skip the pseudocode of the kernel of the second layer as it is similar, but provide the full source code in our code repository. Note that our implementation has multiple points where it could be improved for further performance gains: 1) metadata that is required for the kernels could also be computed with a dedicated kernel to reduce overhead; 2) atomic operations are currently used in the second layer to merge the results from different experts, an alternative implementation that does not use atomic operations could be faster; 3) it could be rewritten in CUDA to make use of dynamic parallelism. We leave those improvements for future work.
In the main paper, we have presented wall-clock time measurements of a single D2DMoE layer. Below, we also ensure that our implementation works and performs well when used for the ViT-B model in which each FFN is replaced with a D2DMoE module. In Figure 11, we measure the averaged processing time and the accuracy of our model. We perform the experiments on an NVIDIA A100 GPU using a batch size of $256$. Each point on the x-axis corresponds to a single $\tau$ threshold and shows the wall-clock time of processing a single input averaged over the entire ImageNet-1k test set. Dynamic inference with D2DMoE offers up to 30% reduction in processing time without affecting the accuracy.
To show that D2DMoE also reduces the execution latency of quantized models, we modify our kernels to handle float16 and int8 data types. In Table 2 we perform a similar experiment to the one from Figure 5. We sample gating decisions from the Bernoulli distribution with probability $p$ and measure the execution time of our experts for the three data type variants.
Table 2: Wall-clock time measurements ($\mu$s) of execution of our D2DMoE layer when using different data types and GPUs.
GPU $\bm{p}$ $\bm{0.0}$ $\bm{0.1}$ $\bm{0.2}$ $\bm{0.3}$ $\bm{0.4}$ $\bm{0.5}$ $\bm{0.6}$ $\bm{0.7}$ $\bm{0.8}$ $\bm{0.9}$ $\bm{1.0}$
RTX 4090 float32 $5$ $9$ $13$ $18$ $23$ $28$ $33$ $38$ $42$ $47$ $52$
float16 $4$ $5$ $7$ $9$ $11$ $14$ $16$ $18$ $21$ $24$ $27$
int8 $4$ $4$ $5$ $7$ $8$ $9$ $10$ $11$ $12$ $13$ $14$
A100 float32 $6$ $9$ $12$ $15$ $19$ $22$ $25$ $28$ $31$ $35$ $38$
float16 $6$ $7$ $8$ $10$ $11$ $13$ $14$ $16$ $17$ $19$ $21$
int8 $7$ $8$ $9$ $10$ $11$ $12$ $14$ $15$ $16$ $17$ $19$
The results show that both the higher activation sparsity (lower $p$) of our method and lower-precision data types are complementary in terms of wall-clock time reduction. While we see a smaller improvement from using int8 over float16 on A100, we attribute this to differences between GPU architectures and software support for low-precision arithmetic.
Appendix D Compatibility with knowledge distillation
Figure 12: Performance of D2DMoE applied on a ViT-S distilled from the larger ViT-B model.
In Section 4.5 we have demonstrated that our method is compatible with two popular model compression methods: quantization and pruning. A natural question is whether our method can be effectively applied to models compressed via knowledge distillation. Since distilled models also exhibit activation sparsity that our method relies on, D2DMoE should be applicable to such models. In Figure 12 we demonstrate the results of D2DMoE when applied on a ViT-S model, which has been trained via knowledge distillation [15] with the torchvision ViT-B being used as the teacher model. We see that D2DMoE is also able to reduce the cost of this smaller model.
Appendix E Routing analysis for large models
Figure 13: Comparision of performance on Gemma-2B for MoEfication with vanilla routing and with our regression routing.
As presented in Figure 4(d), in comparison to other considered benchmarks MoEfication visibly underperforms on language modeling with Gemma-2B. We attribute this to the emergence of massive activations in LLMs that reach a specific scale [40]. Massive activations are outliers along certain feature dimensions whose magnitudes are thousands of times larger than the magnitudes of other activations. The training objective of MoEfication described in Equation 5 uses maximum activation over the entire batch to normalize the target label for each expert. Upon encountering large outlier values, those labels become effectively meaningless, as the values for most of the experts become very close to zero. In this case, the router effectively learns to output zero labels for most of the experts aside from the ones corresponding to the outlier values.
In comparison to MoEfication, our router training scheme does not make use of such normalization, and should therefore be robust to the emergence of massive activations. To validate this, we apply MoEfication on Gemma-2B, but with our regression routing instead of the original router training strategy. We compare the resulting model with vanilla MoEfication in Figure 13 and notice that replacing the routing scheme is enough for the model to learn effective expert assignment, as even though the expert choice is static and the base model is not sparsified, the cost-loss trade-off has significantly improved. This simple experiment shows that our regression routing objective is more robust than MoEfication when scaling to larger models.
Appendix F Router architecture
In comparison to standard linear routers used in MoE models trained from scratch, routers in MoEfication are 2-layer MLPs. To obtain the best performance with D2DMoE, we compare the linear design with MLPs with different hidden sizes for BERT-base and GPT-2-base on Figures 14(a) and 14(b) respectively. Linear routers do not perform well with our method, and overall a 2-layer MLP with a hidden dimension of 128 results in the best performance for both models. Note how for BERT-base, the accuracy curve for a model with the hidden dimension of 128 is slightly worse than for smaller routers, but for harder task with GPT-2 a more complex router is required. Following this analysis, we use 2-layer MLP with a hidden dimension of 128 for most of our experiments in the paper, with the only exception being the larger Gemma-2B model where we scale the hidden dimension accordingly to 512 to match the increase in model dimensionality.
(a) Router architecture ablation for BERT
(b) Router architecture ablation for GPT2
(c) Expert granularity with the GELU.
Figure 14: Additional ablations with router architecture and expert granularity.
Appendix G D2DMoE extension to GLU-based layers
Figure 15: D2DMoE extension to Gated MLP.
To provide better intuition behind the extension of our method to GLU-based gated MLPs mentioned in Section 3.2, we visualize the differences between standard FFN and Gated FFN and the application of our method in Figure 15. Standard Transformer MLP realizes the following function:
$y(x)=\mathbf{W}_{1}A(\mathbf{W}_{2}x),$ (13)
where $\mathbf{W}_{1}$, $\mathbf{W}_{2}$ are the weights for the upscale and downscale projections [^4] We omit biases for simplicity. and $A$ stands for the activation function. In comparison, gated MLP can be written down as:
$y(x)=\mathbf{W}_{1}(A(\mathbf{W}_{g}x)\circ\mathbf{W}_{2}x),$ (14)
where $\mathbf{W}_{g}$ is the weight for the added gate projection.
The intuition behind MoEfication, which our method also follows for standard FFNs, is that the sparsity of the intermediate, post-activation representations determines the sparsity of the output representation. Therefore, the expert split is performed based on the weights of the upscale projection, as zeroed neurons in the upscale activations will also result in zeroed outputs of the downscale projection. When extending D2DMoE to Gated MLPs, our intuition is that the gating projections determine the sparsity of all the later representations, as both upscale and downscale are multiplied with the gating values. Therefore, we propose to build the experts through clustering performed on the gating weights $\mathbf{W}_{g}$ and use the indices obtained through expert split on gating weights to construct experts from $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. Following similar reasoning, for GLU-based models, we also perform activation sparsity enforcement on the gating projections instead of upscale projections as described originally in Section 3.1.
Appendix H Additional results with expert size and GELU
In addition to experiments in Section 5.5, we present the results of similar ablation carried on the sparsified GPT-2 model with GELU activation. The results, presented in Figure 14(c), follow the same pattern as before, which supports our claim that the sparsification enables the GELU-based models to function similarly to ReLU-based ones.
Appendix I Expert activation patterns for attention projection layers
Following the analysis for MoE-converted FFN layers in Section 5.1, we present full results for FFN in Figure 16, and investigate the activation patterns in MHA projections modified with our method in Figures 17, 18, 19 and 20. The projection modules display lower levels of sparsity than FFNs, which is to be expected as our projection layers have lower intermediate dimensionality. Expert selection distribution patterns in $Q$ and $K$ show significant similarity, and the patterns in $V$ and output projections are also similar to a lesser degree. The variance of the number of selected experts in MHA projections is higher than in FFN layers, but it still exists and the distribution in some of the layers seems to be bimodal, which provides further justification for the dynamic-$k$ selection rule.
Appendix J Training and hardware details
In this Section, we describe the technical details used in the D2DMoE conversion procedure. For full reproducibility, we share the source code that we used for conducting the experiments. All experiments were performed using the PyTorch library [29] on the NVIDIA A100 and V100 GPUs on internal clusters. We utilize the fvcore library to count model FLOPs [^5] https://github.com/facebookresearch/fvcore.
J.1 Image classification
All methods start with the same pre-trained ViT-B from the torchvision  [^6] https://pytorch.org/vision/stable/models.html library and are trained on ImageNet-1k using the augmentation proposed by Touvron et al. [44]. We use mixup ($0.8$), cutmix, label smoothing ($0.1$), gradient clipping ($1.0$) and the Adam optimizer with a cosine learning rate schedule without warm-up. For D2DMoE, we replace the MHA projections and train the replacements for $3$ epochs with the initial learning rate $0.001$ and batch size $128$, and then finetune the model for $90$ epochs with sparsity enforcement weight $\alpha=0.2$, initial learning rate $2\cdot 10^{-5}$ and batch size $512$. We then convert the modules into MoE layers, and train the gating networks for $7$ epochs with the initial learning rate set to $0.001$ and batch size $128$. We train ZTW for $100$ epochs in total, allocating $5$ epochs for ensemble training, while keeping the rest of the original hyperparameters unchanged. For MoEfication, we first convert the pre-trained model to ReLU-based one and finetune for $90$ epochs with an initial learning rate of $0.0001$ and batch size $256$. We then split the weights and train the routers for $10$ epochs with the initial learning rate $0.001$ and batch size $256$.
J.2 Text classification
All experiments start from the same pre-trained BERT-base checkpoint. For methods requiring ReLU activation function, we replaced GELU with ReLU and continue model pretraining on concatenated wikipedia [10] and books [55] corpora for $5000$ steps on $8$ GPUs using main setup from https://github.com/huggingface/transformers/blob/main/examples/pytorch/language-modeling/run_mlm.py, per device batch size $96$ and learning rate $5\cdot 10^{-4}$. For MHA projections replacement we use the same corpus and train replaced MLP modules on a single GPU with batch size $128$ and learning rate $0.001$ for $3000$ steps. We finetuned base dense models on CARER dataset for $5$ epochs with $2\cdot 10^{-5}$ learning rate. For sparsity enforcement in D2DMoE we use $\alpha$ linearly increasing from zero to $0.0001$ over training. For both MoEfication and D2DMoE we train routers with batch size 64 and initial learning rate $0.001$ for 5 epochs. In all experiments, we use Adam optimizer with linear learning rate decay. For MoEfication we use expert size $32$, for D2DMoE we use more granular expert size equals $6$. For ZTW we trained ICs for 5 epochs with batch size $32$ and learning rate $0.01$.
J.3 Language modeling
We base our code and hyperparameters for GPT2-base on the nanoGPT repository provided at https://github.com/karpathy/nanoGPT. We initialize the model from https://huggingface.co/openai-community/gpt2. In all pretraining experiments, we initialize models from a publicly available OpenAI checkpoint pre-trained on a closed-source WebText dataset and finetune for the fixed number of 1000 steps with the effective batch size equal to the value in the repository through gradient accumulation. The alpha values for sparsity enforcement can be found at Figure 9(b). We train the routers for D2DMoE and MoEfication for 2000 steps using one GPU and tuning the learning rates for a given expert size from the range between $0.002-0.005$. For router training, we use Adam optimizer and cosine warmup scheduler.
For Gemma-2B, we start from the checkpoint at https://huggingface.co/google/gemma-2b. We also finetune the model for 1k steps with an effective batch size of 1024, sequence length of 1024 and Adam optimizer with a learning rate of 1e-4. As Gemma’s hidden dimension is much larger than the other considered models, we change the hidden dimensionality of the routers to 512 for both our method and MoEfication, but keep the other hyperparameters the same as in the rest of the experiments. For MoEfication, Gemma, we use 512 experts to obtain an expert size comparable to the one in their paper. For our method, we use 2048 experts. In D2DMoE, we set sparsity enforcement weight to $0.00003$. We train the routers for 500 steps with Adam and effective batch size of 16 and use a learning rate of $0.001$.
We report the results for language modeling without the MHA projection replacement step, as we find that it is especially sensitive to changes in the attention layers, which always result in visible loss degradation.
Appendix K Contributions
Filip integrated the codebase and ran the experiments for GPT-2 and Gemma, performed the activation sparsity analysis, and all the analyses on language modeling models. He contributed to the design of dynamic-k gating and played a primary role in designing the experiments and writing the article.
Bartosz set the research direction of the project and proposed the alternative routing scheme, dynamic-k expert selection, the additional activation sparsity enforcement phase for ReLU and GELU, and the replacement of MHA projection layers. He wrote the shared codebase for the experiments, carried out the ViT-B experiments, implemented the custom Triton kernels for the efficient implementation of the method, and also played a primary role in the writing and editing of the article.
Mikołaj made this paper possible by performing all of the experiments at the initial stages of the project and implementing MoEfication and numerous variants of our method. He carried out the BERT experiments, performed weight sparsity compatibility analysis, the ablation study, and contributed to the crafting of the paper.
Simone significantly improved the paper’s readability and provided invaluable advice for revising it.
1def forward_triton_atomic(self, x, routing_tensor):
2 # compute the necessary metadata
3 # split the batch into two groups: executed by that expert or not
4 # (for each expert independently)
5 sort_indices = routing_tensor.argsort(dim=0, descending=True)
6 # get the number of samples executed by each expert
7 expert_bincounts = routing_tensor.sum(dim=0)
8 # actual forward pass
9 intermediate_acts = MoeFirstLayerImplementation.apply(...)
10 final_out = MoeSecondLayerAtomicImplementation.apply(...)
11 return final_out
12
13class MoeFirstLayerImplementation(torch.autograd.Function):
14 @staticmethod
15 def forward(input, weight, bias, sort_indices, expert_bincounts):
16 ...
17 # a grid of kernel instances which divide the computational work
18 # in multiple dimensions: batch dimension (sample_dim),
19 # output dimension (expert_dim) and number of experts dimension (num_experts)
20 grid = (cdiv(sample_dim, BLOCK_SIZE_BD) *
21 cdiv(expert_dim, BLOCK_SIZE_ED), num_experts)
22 moe_first_kernel[grid](...)
23 ...
24
25@triton.jit
26def moe_first_kernel(x_ptr, ...
27 weight_ptr, ...
28 bias_ptr, ...
29 output_ptr, ...
30 sort_indices_ptr, ...
31 expert_bincounts_ptr,
32 ...,
33 ):
34 # based on tl.program_id(axis=0), compute the tile indices
35 # for the batch and output dimensions
36 # (grouped, column major or row major ordering)
37 pid_bd, pid_ed = ...
38 # kernel instances and experts have a many-to-one relationship
39 expert_index = tl.program_id(axis=1)
40 # load the total number of tokens assigned to this expert
41 expert_samples_count = tl.load(expert_bincounts_ptr + expert_index)
42 # calculate the number of instances that need to be used to process all tokens
43 bd_pids_for_expert = tl.cdiv(expert_samples_count, BLOCK_SIZE_BD)
44 # instances that have no computation to perform exit early
45 if pid_bd < bd_pids_for_expert:
46 # calculate offsets that will be used for addressing data in memory
47 offs_bd = ...
48 offs_ed = ...
49 offs_hd = ...
50 # pick the data to load based on the sort indices
51 in_data_indices = tl.load(sort_indices_ptr + expert_index * ... + offs_bd * ...)
52 # calculate memory addresses of the input data and weights
53 # during loading this will group samples for the current learner only
54 x_ptrs = x_ptr + in_data_indices[:, None] * ...
55 w_ptrs = weight_ptr + expert_index * ...
56 # the result will be accumulated in this variable
57 accumulator = tl.zeros((BLOCK_SIZE_BD, BLOCK_SIZE_ED), dtype=tl.float32)
58 # iterate over the innermost dimension
59 for k in range(0, tl.cdiv(hidden_dim, BLOCK_SIZE_HD)):
60 # load the memory from the current input and weight tiles
61 x = tl.load(x_ptrs, mask=..., other=0.0)
62 w = tl.load(w_ptrs, mask=..., other=0.0)
63 # perform matrix multiplication for these tiles and accumulate
64 # (since data is grouped, this can be performed in an efficient manner)
65 accumulator += tl.dot(x, w)
66 # advance the pointers to the next tile
67 x_ptrs += BLOCK_SIZE_HD * stride_x_hd
68 w_ptrs += BLOCK_SIZE_HD * stride_weight_hd
69 # load and add biases to the accumulated result
70 offs_b_ed = ...
71 b_ptrs = bias_ptr + expert_index * ...
72 accumulator += tl.load(b_ptrs, mask=..., other=0.0)
73 # apply the activation function on the result
74 if ACTIVATION == ’relu’:
75 accumulator = relu(accumulator)
76 ...
77 # calculate the memory addresses for the output
78 offs_out_bd = ...
79 out_ptrs = output_ptr + expert_index * ... + \
80 offs_out_bd[:, None] * ... + offs_b_ed[None, :] * ...
81 out_mask = ...
82 # store the result to the main GPU memory
83 tl.store(out_ptrs, accumulator, mask=out_mask)
Listing 1: Simplified pseudocode of our efficient D2DMoE implementation for GPUs
Figure 16: Per-layer distribution of the number of executed experts in D2DMoE trained on the CARER with different $\tau$ thresholds for a standard, non-sparsified model (top row) and a sparsified model (bottom row). The high variability of that number explains the computational gains from using dynamic-$k$.